# Generative AI for Predicting 2D and 3D Wildfire Spread: Beyond Physics-Based Models and Traditional Deep Learning

### v1

@@ 5.4 Wildfire Risk Mapping using Transformers
Wildfire risk mapping plays a critical role in proactive fire management, enabling early warning systems, resource allocation, and strategic response planning across vulnerable landscapes. In recent years, transformer-based deep learning models have emerged as powerful tools for this task, offering scalable and data-efficient solutions for generating accurate and spatially coherent fire risk forecasts.
@@
Limber et al. (2024) developed an intelligent tool to forecast wildfire potential across California by emulating the Wildland Fire Potential Index (WFPI) using a deep learning architecture based on a residual transformer model. The key application area is short-term wildfire risk prediction over large geographic regions, with the goal of improving early warning capabilities for fire management and emergency response. The authors trained the transformer to generate daily WFPI maps using input variables including Daymet meteorological data, MODIS-derived NDVI, and static land fuel classifications from the Scott and Burgan fire behavior fuel models. In this study, the transformer functions as a generative forecasting emulator, producing new spatial fire risk scenarios up to seven days ahead. This model addresses critical challenges in wildfire modeling, including the high computational cost of physical models, and the need for rapid, spatially coherent fire potential forecasts. The residual connection within the transformer leverages temporal autocorrelation in past WFPI values, improving accuracy and stability over short-term forecasts. Notably, the model achieved spatial correlation coefficients ranging from 0.85 to 0.98 across four weekly forecasts in July 2023, demonstrating its effectiveness. The study also incorporated Bayesian hyperparameter optimization to fine-tune the transformer and leveraged high-performance computing resources to efficiently train and evaluate the model. Although no other machine learning models were directly compared, the transformer significantly improves forecast speed and quality relative to existing USGS/USFS WFPI generation methods, highlighting the potential of transformer-based architectures for scalable, data-driven wildfire risk prediction.
6 Visions and Challenges
Generative AI offers a transformative approach to wildfire prediction and management, enabling advanced simulation and analysis of complex fire dynamics. To fully realize its potential, however, several scientific, technical, and operational challenges must be addressed. This section outlines our vision for generative AI in wildfire applications and discusses key barriers to its scalable, trustworthy, and practical deployment.
6.1 New Opportunities with Diverse Generative AI Architecture
Our review finds that current generative AI applications in bushfire prediction mainly rely on GANs, VAEs, and Transformers (Goodfellow et al., 2014; Kingma et al., 2013; Vaswani et al., 2017). However, other architectures—such as Deep AR, Flow-based, Diffusion Models, and Auto-Regressive Transformers—remain underexplored despite their potential to improve wildfire spread prediction, emergency response, and management optimization through more accurate, diverse, and interpretable spatiotemporal scenario generation (Ho et al., 2020; Papamakarios et al., 2021).
Deep AR Models
are a class of generative AI models that generate data sequentially, with each output element conditioned on previously generated elements. This autoregressive property enables them to effectively capture temporal and spatial dependencies in sequential data. In the context of image and spatiotemporal modeling, PixelRNN and PixelCNN are notable examples that extend AR principles to two-dimensional data. PixelRNN employs recurrent neural networks to model the distribution of pixels row by row. They are often used for satellite imagery analysis and Super-resolution (Yang et al., 2019). PixelCNN uses convolutional architectures to achieve similar goals with improved parallelism. Both models generate pixels by conditioning on previously generated pixels, allowing for high-fidelity, structured output. These models have been used in previous studies to generate road, building, precipitation, and land cover maps (Song et al., 2021; Mohammad et al., 2022; Krapu et al., 2024; Wang et al., 2021)
These capabilities make PixelRNN and PixelCNN particularly well-suited for modeling the temporal dynamics and spatial evolution of wildfire progression, where the state at any given time or location depends heavily on prior conditions (Van Den Oord et al., 2016).
Flow-Based Models
are a class of generative AI models that learn invertible transformations between simple base distributions (e.g., Gaussian) and complex data distributions through a series of bijective mappings. Unlike other generative AI models, flow-based models offer exact and tractable likelihood computation, making them particularly attractive for scientific applications requiring probabilistic interpretability. Notable examples include RealNVP (Dinh et al., 2017) and Glow (Kingma and Dhariwal, 2018), which use affine coupling layers and invertible 1x1 convolutions to enable efficient sampling and density estimation. Their ability to model high-dimensional, continuous variables with fine-grained control makes them especially well-suited for simulating geospatial variables such as fuel moisture content, temperature gradients, or wind vector fields—factors that are critical in wildfire behavior prediction. These models can learn complex, multivariate relationships directly from data while preserving physical consistency due to their invertibility and structure-preserving transformations.
Diffusion Models
are a class of generative AI models that learn to generate data by reversing a gradual noising process. During training, data is iteratively corrupted by adding Gaussian noise, and the model learns to reverse this process step-by-step to recover the original input. This denoising diffusion process enables the model to capture complex data distributions with high fidelity. Notable examples include the Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020), which has demonstrated state-of-the-art performance in generating high-resolution, diverse images. Diffusion models are particularly well-suited for reconstructing missing data and generating plausible future environmental scenarios from noisy or sparse 2D inputs, such as satellite imagery or geospatial raster data. Their ability to progressively refine samples allows them to infer fine-grained spatial patterns and fill in incomplete observations with high accuracy, making them powerful tools for environmental modeling tasks like wildfire risk mapping or weather variable simulation.
Each generative architecture contributes distinct strengths to wildfire modeling. Autoregressive (AR) models are particularly effective for capturing temporal dependencies, making them well-suited for forecasting the sequential dynamics of fire movement over time (Limber et al., 2024). Flow-Based Models offer invertible mappings that preserve exact likelihoods, allowing them to model spatially continuous variables with high fidelity—ideal for representing terrain, fuel distribution, or meteorological gradients (Dinh et al., 2016; Kingma and Dhariwal, 2018). Diffusion Models, which have gained recent popularity, iteratively denoise sampled data to synthesize high-quality outputs such as 2D raster maps or geo-referenced images, enabling scenario-based simulation of fire evolution under diverse environmental conditions (Ho et al., 2020; Rombach et al., 2022). Finally, Auto-Regressive Transformers support the fusion and contextual learning of multiscale, multimodal data—such as weather, vegetation, topography, and sensor inputs, facilitating comprehensive, high-resolution prediction and decision support for wildfire risk assessment and management (Vaswani et al., 2017; Xu et al., 2024b).
@@ 6.1 New Opportunities with Diverse Generative AI Architecture
6.2 Visions for Generative AI-powered Wildfire Applications
This section outlines a series of forward-looking visions for how generative AI, particularly in the form of foundation models, multimodal transformers, diffusion models, and LLM-based agents, can be harnessed to reshape wildfire prediction and management across scales and settings (as illustrated in Figure 5).
6.2.1 Advancing Multimodal Approaches in 2D and 3D Wildfire Modeling
@@
Traditional wildfire modeling has primarily relied on 2D spatial datasets—such as satellite imagery, temperature maps, and vegetation indices. However, fire behavior is inherently three-dimensional, shaped by terrain elevation, vertical fuel distribution, atmospheric conditions, and urban infrastructure. Conventional deep learning approaches often handle 2D and 3D data in isolation, requiring separate models for different modalities, which limits integration and hampers model adaptability. The emergence of high-fidelity 3D fire spread modeling offers a transformative opportunity to improve emergency response, evacuation planning, and long-term mitigation (Moinuddin et al., 2024). By accounting for the vertical complexity of fire dynamics, 3D models enhance the accuracy of predictions related to fire pathways, temperature distribution, and ember dispersion, while also supporting immersive visualizations (Xu et al., 2023a).
Recent advancements in generative AI architectures, particularly VAEs and transformers, now allow for the development of unified multimodal models capable of processing geospatial, meteorological, and 3D structural data within a single framework. These models leverage shared latent representations and cross-modal attention to learn spatial, temporal, and structural patterns holistically. Unlike conventional systems that rely on fusing outputs from separate networks, this end-to-end generative approach improves maintainability, scalability, and adaptability. By embracing generative multimodal modeling, researchers and practitioners can move beyond fragmented workflows toward more cohesive, real-time predictive systems.
6.2.2 Developing a Conversational Multi-Agent AI Assistant for Wildfire Prediction Using LLMs and RAG
The integration of Large Language Models (LLMs) with Retrieval-Augmented Generation (RAG) offers a novel pathway for developing conversational multi-agent AI assistants for wildfire prediction. These systems can engage in multi-turn dialogue, retrieve domain-specific wildfire data, and generate context-aware, user-adaptive insights. Building on successful use cases in transportation and climate science (Xu et al., 2024e; Xie et al., 2025), this approach shifts wildfire intelligence from static interfaces to interactive, human-centered decision support. In this framework, specialized agents handle sub-tasks such as fire spread simulation, atmospheric analysis, or evacuation planning, coordinated by a dialogue manager driven by a fine-tuned LLM. RAG ensures the assistant’s responses are not only generative but grounded in real-time data and scientific knowledge, drawing from sources like localized fire weather indices or historical ignition patterns (Xie et al., 2025).
This architecture enables natural language interaction for urban planners, emergency responders, and community members through an intelligent AI assistant capable of automatically retrieving relevant data, simulation outputs, and optimized solutions for decision support. It presents this information in an accessible, easy-to-understand format, providing non-technical users with tailored recommendations and real-time updates. Unlike traditional approaches that require manual data collection, technical expertise, and simulation execution, often resulting in outputs filled with domain-specific jargon that hinder interdisciplinary collaboration, the AI assistant offers a unified, scalable, and explainable interface. Continuously learning from expert feedback and new data streams, it supports adaptive wildfire intelligence. When embedded in smart city systems, such assistants foster more proactive, transparent, and inclusive fire risk assessment and emergency response.
6.2.3 Building an AI Foundation Model for Interdisciplinary Wildfire Management
The emergence of AI foundation models—large-scale, pre-trained systems built primarily on transformer architectures—has opened new possibilities for addressing complex, interdisciplinary challenges in environmental and urban domains (Li et al., 2024a; Myers et al., 2024). By nature, an AI foundation model is a large-scale, general-purpose model trained on diverse data to support a wide range of downstream tasks, serving as a flexible backbone for AI systems—not limited to any specific machine learning or deep learning technique (Li et al., 2024a). These models enable cross-domain knowledge transfer, zero-shot inference, and rapid fine-tuning, making them well-suited for high-impact applications such as wildfire prediction. A foundation model for wildfire management would leverage generative AI architectures, including LLMs and multimodal transformers, to learn from diverse datasets across spatial, temporal, and semantic dimensions. Trained on inputs such as satellite imagery, weather records, terrain data, fuel assessments, and fire incident reports, the model could generate predictive outputs and scenario-based simulations under varying conditions.
Unlike traditional models or task-specific deep learning pipelines, a generative foundation model offers a unified, adaptive, and scalable framework. Recent advancements in generative AI architectures—such as transformers, diffusion models, and autoregressive networks—have paved the way for developing multimodal AI models that seamlessly integrate diverse data sources, including satellite imagery, LiDAR, weather time series, textual reports, and both 2D and 3D spatial data. These models excel at capturing cross-modal relationships, enabling robust generalization across varied fire scenarios, geographic regions, and management tasks. Unlike traditional deep learning methods, such as CNN-based models specialized for 2D raster GIS data, AI foundation models for wildfire management harness the multimodal and multipurpose capabilities of contemporary generative architectures. This allows for the incorporation of evolutionary features utilizing both 2D and 3D data, along with other modalities, to not only enhance the accuracy of fire risk predictions across multiple dimensions but also support critical wildfire management tasks, such as detecting smoke, fire, and vulnerable vehicle and infrastructure from CCTV feeds, within a unified AI framework. Such an approach enhances the deployability, scalability, and maintainability of AI-powered systems in real-world applications.
6.2.4 Computationally Efficient Hypothetical Wildfire Scenario Generation on Edge Devices
@@
Generative AI models such as VAEs and Transformers have emerged as computationally efficient surrogates for simulating wildfire spread and fire propagation. Unlike physics-based models that rely on intensive numerical solvers, these generative architectures can emulate spatiotemporal fire dynamics from historical or synthetic data, enabling the rapid generation of high-resolution wildfire scenarios with minimal computational overhead. Once trained, they deliver near-instant predictions, making them ideal for time-critical tasks such as emergency response, evacuation planning, and real-time firefighting operations. A key advantage of these models is their ability to run on edge devices—including UAVs, remote sensors, connected vehicles, and mobile platforms like tablets or smartphones used by firefighters. Prior work has shown that convolutional and transformer-based models can be optimized for execution on low-power hardware such as NVIDIA Jetson Nano or Intel Movidius, enabling on-board inference without cloud dependence (Hu et al., 2024; Mahdi and Mahmood, 2022). This reduces reliance on remote infrastructure and mitigates communication delays or outages, especially in remote or disaster-affected areas (Xu et al., 2024d).
@@
These edge-deployable models also support decentralized computing, enabling scenario generation using the most updated real-time, field-collected data such as wind speed, terrain, vegetation, and humidity. Lightweight Vision Transformers (ViTs) and compact CNN-based fire segmentation algorithms are increasingly compatible with embedded systems, supporting high-frequency updates and real-time forecasting on resource-constrained devices (Spiller et al., 2022; Lee et al., 2024). This edge-oriented paradigm empowers frontline responders with on-the-fly fire modeling, delivering predictive fire trajectories, containment strategies, and dynamic risk maps—all processed locally. Embedded within mobile edge computing frameworks, these systems can continuously adapt to incoming sensor data, offering scalable, resilient, and connectivity-independent wildfire intelligence.
6.2.5 Interactive Wildfire Management through Explainable AI and Cognitive Digital Twins
Urban digital twins are dynamic virtual representations of physical city systems that integrate real-time data, simulations, and models to support planning, monitoring, and decision-making (Xu et al., 2024c; Weil et al., 2023). These digital twin technologies have been widely adopted across various smart city domains, including intelligent transportation systems, smart buildings, and environmental monitoring and management (Xu et al., 2023b; li2020towards; Diakite et al., 2022). Building upon this foundation, cognitive digital twins represent a new paradigm by embedding AI into digital twin frameworks to enhance their ability to learn, reason, and adapt (Zheng et al., 2022). Unlike conventional twins, cognitive digital twins integrate human-in-the-loop mechanisms to improve the trustworthiness, transparency, and interpretability of the decisions and predictions derived from underlying data models and simulations (Niloofar et al., 2023).
A key advantage of developing cognitive digital twins powered by generative AI is their ability to support explainable AI approaches, enabling these systems—often deployed as smart city apps or web platforms—to interact effectively with human users and enhance the supervision and interpretability of AI-generated results. Many generative AI models, such as VAEs and Transformers, rely on latent spaces and latent representations to encode complex data structures into compressed, abstract forms. These latent variables can be visualized and interpreted to reveal how models process information and generate predictions or classifications (Xu et al., 2024a; Liu et al., 2019). This capability is especially valuable in cognitive digital twin systems, where enhanced model explainability and interpretability are essential for ensuring accountability and transparency—particularly in time-critical emergency response scenarios such as wildfire spread prediction and risk forecasting (Abdollahi and Pradhan, 2023). In such applications, the multivariate feature space of complex environmental datasets can be encoded into a lower-dimensional latent space, enabling visualization and the extraction of meaningful patterns (Liu et al., 2019). Through interactive visualization and visual analytics, these insights can support a range of downstream tasks, including anomaly detection, decision support, and multivariate clustering for risk analysis and fire event characterization (Kwon et al., 2023; Xu et al., 2024a), and are essential for increasing the trustworthy of the prediction outcomes from generative AI models.
Meanwhile, the cyberinfrastructure supporting a digital twin can serve as an automated pipeline for integrating real-time environmental data from various sources, including sensors, unmanned aerial vehicles (UAVs), and crowdsourcing platforms. It also provides a cloud-based computing environment capable of hosting generative AI models developed for wildfire prediction. In this setting, diverse environmental data streams can be seamlessly ingested and processed by the AI models, ensuring that wildfire forecasts are generated using the most up-to-date information available.

### v2

@@ 5.3 Wildfire Risk Mapping
Wildfire risk mapping plays a critical role in proactive fire management, enabling early warning systems, resource allocation, and strategic response planning across vulnerable landscapes. In this review, only one study using transformer-based method have been discovered.
@@
Limber et al. (2024) developed a residual transformer model to forecast wildfire potential across California by emulating the Wildland Fire Potential Index (WFPI), aiming to enhance short-term risk prediction and early warning for fire management. The model was trained to generate daily WFPI maps using Daymet meteorological data, MODIS-derived NDVI, and static fuel classifications from the Scott and Burgan fire models. As a generative forecasting emulator, it produces new spatial fire risk scenarios up to seven days ahead, addressing the high computational demands of physical models while ensuring rapid, spatially coherent forecasts. The residual connection leverages temporal autocorrelation in WFPI to boost short-term accuracy and stability.
As a performance summary, the model achieved spatial correlation coefficients of 0.85–0.98 across four weekly forecasts in July 2023, showing a slight tendency to underpredict extreme risk values. Precision remained high for one- and two-day forecasts, with modest degradation from days three to seven due to error propagation. Bayesian hyperparameter optimization and HPC resources enabled efficient tuning and training—requiring 42 hours on 48 GPUs for tuning and 24 hours on 24 GPUs for training—while full inference for four weeks of statewide daily forecasts was completed in just 6.5 minutes on a 128-core CPU. Although no direct comparisons were made to traditional or other ML models, the transformer substantially improved forecast speed and quality over existing USGS/USFS WFPI approaches, underscoring its promise for large-scale, data-driven wildfire risk prediction.
6 Discussion and Future Directions
Our targeted literature review identified and critically examined 11 studies that specifically explore the application of generative AI models to wildfire science. This relatively small number underscores that the integration of generative methods in this domain remains nascent, with substantial opportunities for further research and methodological innovation. Among the works reviewed, the predominant focus has been on leveraging GAN, VAE, and Transformer architectures to enhance bushfire prediction capabilities Goodfellow et al. (2014); Kingma et al. (2013); Vaswani et al. (2017). This clear research gap highlights a fertile area for advancing wildfire science through the development of more sophisticated, data-driven generative approaches that can simulate complex spatiotemporal fire dynamics and support robust decision-making.
Building on the preceding review of current generative AI applications in wildfire modeling and prediction, our analysis of 11 studies shows that generative AI models can achieve accuracies around 90%, with some demonstrating computational performance exceeding traditional methods by several orders of magnitude. Combined with recent trends in LLM-powered agentic AI and distributed edge computing, these advantages open promising new research directions that remain largely unexplored. This section synthesizes key insights from the literature, outlines future avenues for investigation, and highlights major challenges along with potential strategies to address them.
@@ 6 Discussion and Future Directions
6.1 Future Research Directions for Generative AI-powered Wildfire Applications
This section outlines a series of forward-looking research directions for how the existing generative AI-powered fire spread prediction applications can be further extended with more resaerch and development efforts to reshape wildfire prediction and management across scales and settings (as illustrated in Figure 3).
6.1.1 Unified Simulation Framework for 2D and 3D Wildfire Dynamic
@@
Previous studies have highlighted the importance of using multimodal data to improve bushfire prediction, while also noting the significant technical challenges involved in fusing heterogeneous datasets Shadrin et al. (2024); Zakari et al. (2025). Traditional wildfire modeling has primarily relied on 2D spatial inputs—such as satellite imagery, temperature maps, and vegetation indices—even though fire behavior is inherently three-dimensional, shaped by terrain elevation, vertical fuel structures, atmospheric dynamics, and built environments. Conventional deep learning approaches typically handle 2D and 3D data in isolation, requiring separate models for different modalities, which limits integration and adaptability.
Despite the experimental use of transformers and VAEs for fire spread prediction in the 11 reviewed studies, current research has not explored their potential as multimodal generative frameworks for unifying diverse data sources. From a data science standpoint, architectures like VAEs and Transformers are inherently well-suited to integrate heterogeneous inputs by learning joint representations across modalities. VAEs can encode geospatial, meteorological, and 3D structural information into shared latent spaces governed by probabilistic distributions, while Transformers leverage self-attention to dynamically capture cross-modal and long-range dependencies Li et al. (2025b); Xu et al. (2023). This capability removes the need for explicit feature concatenation or separate fusion pipelines, enhancing scalability and allowing models to holistically capture the complex interplay of factors that drive wildfire propagation. By jointly learning how dynamic weather patterns interact with terrain and fuel characteristics, such approaches could improve predictive accuracy and enable scenario-based forecasting under varying meteorological conditions—critical for informed emergency response planning and efficient resource allocation. These considerations highlight a promising research agenda that leverages the unique strengths of multimodal generative AI to advance holistic wildfire simulation.
6.1.2 Chatbots for Wildfire Decision Intelligence
Recent trends across various scientific disciplines, including transportation and climate science, increasingly leverage the integration of Large Language Models (LLMs) with Retrieval-Augmented Generation (RAG) to develop advanced conversational AI systems Xu et al. ; Xie et al. (2025). These multi-agent frameworks enable interactive, multi-turn dialogue that dynamically retrieves domain-specific data and generates context-aware, user-adaptive insights. Despite this growing popularity in other fields, such paradigms have yet to be explored for supporting wildfire decision intelligence.
This represents a compelling research direction. Future work could explore how conversational AI assistants—built on fine-tuned LLMs, RAG mechanisms, and an agentic AI paradigm—might transform wildfire prediction and management Xu et al. ; Xu et al. (2025). In such a framework, specialized agents could handle sub-tasks such as fire spread simulation, atmospheric analysis, or evacuation planning, all coordinated by a dialogue manager to ensure coherent, user-centered interactions. RAG would ground the assistant’s responses in real-time wildfire data and scientific knowledge, drawing on sources like localized fire weather indices and historical ignition patterns.
By enabling natural language interaction, these AI assistants could empower urban planners, emergency responders, and community members to access relevant data, simulation outputs, and optimized decision solutions in an accessible format. This contrasts sharply with traditional systems that often rely on manual data processing and technical interfaces dense with domain-specific jargon. As an adaptive platform that learns continuously from expert feedback and evolving data streams, a wildfire-focused conversational AI system could facilitate more proactive, transparent, and inclusive approaches to fire risk assessment and emergency response. Advancing this vision represents an important future research agenda, leveraging the strengths of emerging LLM-RAG architectures to address longstanding challenges in wildfire decision support.
6.1.3 AI Foundation Model for Interdisciplinary Wildfire Management
The emergence of AI foundation models—large-scale, pre-trained systems typically built on transformer architectures—has become an increasingly popular trend across numerous scientific domains, including remote sensing, urban logistics optimization, and GIS Li et al. (2024a); Myers et al. (2024). These models serve as general-purpose backbones trained on diverse datasets to support a wide range of downstream tasks, enabling cross-domain knowledge transfer, zero-shot inference, and rapid fine-tuning. Despite their growing success in other disciplines, foundation models remain largely unexplored in the context of wildfire science.
This gap represents a promising direction for future research. A foundation model for wildfire management could leverage generative AI architectures—such as LLMs, multimodal transformers, or diffusion networks—to integrate heterogeneous datasets across spatial, temporal, and semantic dimensions. Trained on inputs like satellite imagery, LiDAR scans, weather time series, terrain maps, fuel assessments, and incident reports, such a model could generate predictive outputs and scenario-based simulations under varying conditions, supporting applications ranging from fire spread forecasting to resource deployment planning.
Unlike traditional task-specific deep learning pipelines that often rely on separate models for different data modalities—such as CNNs for 2D raster GIS data—an AI foundation model would offer a unified, adaptive, and scalable framework. By inherently capturing cross-modal relationships, such a model could robustly generalize across diverse geographic regions, fire regimes, and management objectives. Integrating multimodal data within a single generative architecture would enable simultaneous analysis of evolving 2D and 3D spatial features, detection of smoke and vulnerable infrastructure from CCTV feeds, assessment of wildfire impacts on wildlife, and even prediction of housing price shifts in fire-prone areas, alongside automated synthesis of textual reports and advisories. Advancing this line of research represents a significant opportunity to enhance the deployability, scalability, and maintainability of AI-powered wildfire decision support, establishing a transformative paradigm for interdisciplinary wildfire management.
6.1.4 Real-time Fire Scenario Generation on Mobile Devices
@@
Generative AI models such as VAEs and Transformers have shown considerable promise as computationally efficient surrogates for simulating wildfire spread. Unlike physics-based models that rely on intensive numerical solvers, these architectures can emulate spatiotemporal fire dynamics from historical or synthetic data, producing high-resolution wildfire scenarios with minimal computational overhead. Once trained, they offer near-instant inference, making them attractive for time-critical applications such as emergency response, evacuation planning, and real-time firefighting operations, and potentially deployable on mobile devices with limited computing resources.
However, despite the reviewed 11 studies demonstrating the computational efficiency and superior inference speeds of generative AI models, this research remains largely at an experimental stage. None of these studies have explored the bright potential of deploying such lightweight generative models directly on mobile platforms and edge devices. This represents an important gap and a promising direction for future research.
Edge and mobile deployment could distribute computational and data analytics workloads across devices such as smartphones, UAVs, sensor controllers, and CCTV camera systems, reducing reliance on remote infrastructure and mitigating communication delays—especially critical in disaster-affected or connectivity-limited regions. Prior work has shown that convolutional and transformer-based models can be optimized for low-power hardware, such as NVIDIA Jetson Nano or Intel Movidius, enabling onboard inference without cloud dependence Hu et al. (2024); Mahdi and Mahmood (2022). Lightweight ViTs-based fire segmentation models are increasingly compatible with embedded systems, supporting high-frequency updates and real-time forecasting on resource-constrained devices Spiller et al. (2022); Lee et al. (2024).
@@
Future work could leverage these developments to embed generative wildfire models within mobile edge computing frameworks, enabling on-the-fly scenario generation using the most current field-collected data—such as wind speed, terrain, vegetation, and humidity. This would empower frontline responders with localized fire trajectory predictions, dynamic risk maps, and adaptive containment strategies, all processed directly on-site. Advancing this research agenda could establish a scalable, resilient, and connectivity-independent paradigm for delivering real-time wildfire intelligence.
6.1.5 Trustworthy Wildfire Prediction via Explainable AI and Interactive Visual Analytics
Recent advances in generative AI have opened new avenues for developing explainable and interactive wildfire prediction systems that go beyond traditional black-box models. Many state-of-the-art generative architectures, such as VAEs and Transformers, rely on latent spaces to encode complex environmental data into compressed representations. These latent variables offer a powerful mechanism for uncovering how models process input features and arrive at predictions or classifications Xu et al. (2024a); Liu et al. (2019). Notably, although the 11 reviewed studies employed VAE- and transformer-based approaches that inherently learn latent representations and embeddings, none explored explainable AI techniques that visualize or interpret these latent spaces to clarify the models’ decision-making processes.
This gap highlights a compelling direction for future research. By developing visual analytics interfaces that directly examine latent spaces, stakeholders—including emergency managers, urban planners, and community members—could better understand how inputs such as fuel conditions, topography, and weather patterns collectively influence predicted fire spread. Interactive exploration of these low-dimensional representations would facilitate detection of meaningful patterns, anomalies, or risk clusters, enabling non-technical users to interrogate AI-driven forecasts and fostering greater transparency and trust. Such approaches stand in contrast to existing wildfire modeling tools that often produce opaque outputs requiring specialized interpretation.
While these explainable AI systems could be embedded within broader platforms—such as urban digital twins or smart city dashboards—the core research agenda centers on advancing methodologies for visualizing, interpreting, and interacting with the internal decision processes of generative AI wildfire models. Integrating real-time data streams from IoT sensors, UAVs, or crowdsourcing platforms, these systems could continuously update their analyses and present the latest insights in an intuitive, user-centered format. Pursuing this line of research promises to significantly enhance the interpretability, reliability, and adoption of AI-based wildfire prediction tools, paving the way for more transparent and participatory fire risk management.